OpenStudy (anonymous):

are 2x=10 and 6+x=7 linear equations?

Directrix (directrix):

They are vertical lines which are linear equations. 2x=10 simplifies to x = 5. What does this one simplify to: 6+x=7 @Georgia_peach Note: Vertical lines do not have slope, or, as some say, vertical lines have undefined slope.

OpenStudy (agent0smith):

Linear equations only have multiplication or division of variables by constants, as well as addition and subtraction. eg. 4x-y=0, y=7x-2, y/5+x/7=1 if you see something like x^2, or y^2, or sqrt(x), or xy or yx etc... the equation is NOT linear

OpenStudy (agent0smith):

You also can not have an x or y in a denominator, like 1/y or 2/x would make the equation not linear.

OpenStudy (anonymous):

kk... Thanks yall, I understand it pretty much better now.

Directrix (directrix):

@agent0smith and I were happy to assist you.

OpenStudy (anonymous):

y = 3x - 6 So if I'm right this is a equation

OpenStudy (agent0smith):

it's a LINEAR equation, yes.

OpenStudy (agent0smith):

An equation just needs an equals sign.